OpenStudy (anonymous):

medal and fan awarded what is the exact value of the expression sqrt4y^8 + sqrty^2 - 2sqrty^2 + sqrty^8? simplify is possible 5y^4 + 3y 7y^4 3y^4 + 4y 7y^6

OpenStudy (rajee_sam):

Do you know how to take √ of numbers and √ of exponents?

OpenStudy (anonymous):

no.

OpenStudy (rajee_sam):

now consider the first term\[\sqrt{4y ^{8}}\]

OpenStudy (rajee_sam):

Now this can split as \[\sqrt{4} .\sqrt{y ^{8}}\]

OpenStudy (anonymous):

ok

OpenStudy (rajee_sam):

now what is √ 4 ?

OpenStudy (anonymous):

2

OpenStudy (rajee_sam):

good. Now for taking the sqrt of exponents. Note: You have keep the exponents in even nos. now you have y^8 so leave it like that. if you have y^3 split it as y^2.y Now we have y^8. Taking the √ of exponents is just dividing the exponent by 2, meaning dividing 8/2 and writing y with the new exponent. so \[\sqrt{y ^{8}} = y ^{4}\]

OpenStudy (anonymous):

ok

OpenStudy (rajee_sam):

so now you first term reduced to \[2y ^{4}\]

OpenStudy (rajee_sam):

Now can you do the second one. you have √ y² . what will be the sqrt of this?

OpenStudy (rajee_sam):

divide the power by 2 and write the letter with the new exponent.

OpenStudy (rajee_sam):

should not take this much time

OpenStudy (rajee_sam):

gtg can't wait the whole day for you to answer